# UniHR: Hierarchical Representation Learning for Unified Knowledge Graph Link Prediction

### v5

@@ Appendix B A Decoder Analysis
To explore the effectiveness of our UniHR encoding further, we pair UniHR with different decoders and evaluated them on triple prediction task. In addition to the previously mentioned unified framework UniHR + Transformer, we also experiment on UniHR + ConvE with two scoring strategies. The ConvE (Dettmers et al. 2018) is the decoder customized for triples and its scoring function is $vec\left(\sigma\left(\left[\mathbf{\tilde{h}}_{h};\mathbf{\tilde{e}}_{r}\right]*\psi\right)\right)$, where $\mathbf{\tilde{h}}_{h}$ and $\mathbf{\tilde{e}}_{r}$ represent reshaped 2D embeddings of head entity $h$ and relation $r$, and * is a convolution operator. The $vec\left(\cdot\right)$ and $\psi$ are denoted as the vectorization function and a set of convolution kernels.
@@
Table 6: Base link prediction on FBHE, FBH and DBHE. All baselines’ results are taken from (Xiong et al. 2024). The best results among all models are written bold, while the second are underlined. The $s_{f}$ and $s_{h}$ denote $\left(f,has\;head\;entity,h\right)\ \left(f,has\;tail\;entity,t\right)$ and $\left(h,r,t\right)$ two types of scoring method respectively.
Due to our special representation, there exists two scoring methods for atomic triples, thus we present the base link prediction results separately for each scoring method. The $s_{f}$ represents scoring triples $\left(f,has\;head\;entity,h\right)\$and $\left(f,has\;tail\;entity,t\right)$, and $s_{t}$ represents scoring $\left(h,r,t\right)$. The performance of base link prediction is shown in Table 6. It can be observed that regardless of the scoring method employed, we both achieve competitive performance, especially with scoring $\left(h,r,t\right)$ on FBH and scoring $\left(f,has\;head\;entity,h\right)\,\left(f,has\;tail\;entity,t\right)$ on DBHE. We attribute the differences in performance under different scoring methods to dataset characteristics. The DBHE dataset is relatively smaller, and scoring method $s_{f}$ effectively alleviates overfitting problem. Conversely, for larger datasets FBH, scoring based on $\left(h,r,t\right)$ minimizes information loss.
Table 8 shows the results of triple prediction on three benchmark datasets. Among all baselines, Quate and Bique struggle to model the mapping relationship between atomic facts and nested facts. Furthermore, prior works (Chung and Whang 2023) do not guarantee that all atomic facts in the nested fact test set are present in the training set as entities, which shifts the problem from a transductive setting to an inductive setting, leading to significant performance gaps between these baselines. On most metrics, our method outperforms BiVE and NestE which are specifically modeled for nested facts. Notably, NestE fully preserves the semantics of atomic facts. However, on the FBHE dataset, UniHR + ConvE achieves an improvement of 0.58 (6.4%) points in MRR and 0.24 (2.4%) points in Hits@10 compared to the state-of-the-art model NestE and the second-best performance after UniHR + Transformer on the FBH and DBHE datasets, demonstrating UniHR’s powerful graph structure encoding capabilities. We also carry out ablation experiments on UniHR + ConvE as shown in Table 8. Performance declines are observed after removing any part of the HiSL module, showing the significance of HiSL for hierarchical encoding.
Variant FBH DBHE
MR MRR H@10 MR MRR H@10
w/o initial $\textbf{h}_{f}$ 5.22 0.909 0.980 2.56 0.794 0.978
w/o $\textbf{W}_{r}$ 3.06 0.944 0.992 2.30 0.850 0.976
w/o intra-fact MP 3.97 0.897 0.972 2.02 0.842 0.983
w/o $\omega_{\tau(r)}$ 2.70 0.934 0.992 2.69 0.810 0.973
w/o $\textbf{W}_{\lambda(r)}$ 2.50 0.941 0.992 2.37 0.810 0.978
w/o inter-fact MP 2.61 0.913 0.991 2.11 0.827 0.986
UniHR 2.46 0.946 0.993 1.90 0.862 0.987
Table 7: Ablation studies on triple prediction task.
@@
Table 8: Triple prediction on FBHE, FBH and DBHE. All baselines’ results are taken from (Xiong et al. 2024). The best results among all models are written bold. For BiVE (Chung and Whang 2023) and NestE (Xiong et al. 2024), we pick their variants with best performance.
Appendix C B More Ablation Study
We present more ablation experiments on triple prediction. As shown in Table 7, all variants with certain modules or parameters removed exhibit a decrease in performance. We can conclude that intra-fact and inter-fact message passing modules both play crucial roles, allowing UniHR to more fully represent the current fact node with enhanced structural information. We also change the initialization of fact node embeddings to learnable embeddings (w/o initial $\textbf{h}_{f}$), and results indicate that initializing fact node representations is essential. It highlights key information in the facts, mitigating the noise introduced by excessive neighbors.
@@ Appendix C B More Ablation Study
GNN_intra-fact dropout 0.1 0.1 0.1 0.3 0.1 0.3 0.2 0.1 0.1
GNN_inter-fact activation tanh tanh tanh tanh tanh tanh tanh tanh tanh
GNN_dropout 0.1 0.1 0.1 0.3 0.1 0.3 0.2 0.1 0.1
@@
Appendix D C More Related Works
@@
Most existing techniques in KG representation learning are proposed for triple-based KGs. Among these techniques, knowledge graph embedding (KGE) models (Bordes et al. 2013; Sun et al. 2018) have received extensive attention due to their effectiveness and simplicity. The idea is to project entities and relations in the KG to low-dimensional vector spaces, utilizing KGE scoring functions to measure the plausibility of triples in the embedding space. Typical methods include TransE (Bordes et al. 2013), RotatE (Sun et al. 2018), and ConvE (Dettmers et al. 2018).
Depending on the KGE model alone has limitation of capturing complex graph structures, whereas augmenting global structural information with a graph neural network (GNN) (Vashishth et al. 2019; Nathani et al. 2019; Xu, Bao, and Liu 2023) proves to be an effective approach for enhancement. The paradigm of combining GNN as encoder with KGE scoring function as decoder helps to enhance the performance of KGE scoring function. These GNN methods design elaborate message passing mechanisms to capture the global structural features. Typically, CompGCN (Vashishth et al. 2019) aggregates the joint embedding of entities and relations in the neighborhood via a parameter-efficient way and MA-GNN (Xu, Bao, and Liu 2023) learns global-local structural information based on multi-attention. These methods achieve impressive results on triple-based KGs but are challenging to generalize to beyond-triple KGs. Additionally, these GNNs usually focus more on modeling global information while neglecting local information.
@@
Firstly, real-world data sources naturally contain heterogeneous data forms (Liu et al. 2025b). For example, Wikidata (Vrandečić and Krötzsch 2014) includes triple-based facts, hyper-relational facts, temporal facts, and even nested facts, financial reports and research papers usually contain various modalities such as tables, images, and text. Unified representation aims to integrate multiple heterogeneous data within a single semantic structure, avoiding the need for cross-type data conversion or independent processing during use. In many real-world scenarios (e.g., question answering, recommendation, or transfer learning), the research community is constantly exploring the benefits of unified representations. For question answering (QA), TrustUQA (Zhang et al. 2025b) demonstrates a unified representation allows the knowledge base to contain a broader range of knowledge types, thereby simplifying the burden of the knowledge retrieval module and enabling the QA system to cover a wider variety of questions and answers. For recommendation systems, DURation (Lu et al. 2022) with unified representation effectively eliminate information redundancy caused by separate module designs and help capture cross-dimensional relational knowledge, thus improving the comprehensiveness and accuracy of recommendations. Moreover, UniHR not only meets the requirements of real-world KG reasoning tasks (involving multiple types of facts), but we also believe that unified representation itself is beneficial for transfer learning. It not only facilitates transfer learning across different types of KGs, but also enables pre-training across multiple KG types (as demonstrated in Section 5.4). This further promotes more effective pre-training and transfer learning in the field of knowledge graphs.
Appendix E D Detailed Results of Our “Potential”
@@
Appendix F E Dataset Details
@@
Among them, WikiPeople is a dataset derived from Wikidata (Vrandečić and Krötzsch 2014) concerning entities type “human”. WikiPeople filter out the elements which have at least 30 mentions as key-value pairs. WD50K is a high-quality dataset extracting from Wikidata statements and avoiding the potential data leakage which allows triple-based models to memorize main fact in the H-Facts of test set. The “with Q(%)” column in Table 9 denote the number of facts with auxiliary key-value pairs and the “Arity” column denote range of the number of entities in hyper-relational facts.
Nested factual knowledge graph datasets FBH and FBHE (Chung and Whang 2023) are constructed based on FB15k237 from Freebase (Bollacker et al. 2008) while DBHE is based on DB15K from DBpedia (Lehmann et al. 2015). FBH contains nested facts that can be only inferred inside the atomic facts, while FBHE and DBHE contain externally-sourced nested relation crawling from Wikipedia articles, e.g., NextAlmaMater and SucceededBy.
Temporal knowledge graph dataset wikidata12K is also a subset of Wikidata (Vrandečić and Krötzsch 2014), which represents the time information $\tau\in\mathcal{T}$ as time intervals.
Appendix G F Hyperparameter Settings

### v6

@@ Conclusion
Acknowledgements
This work is founded by National Natural Science Foundation of China (NSFC62306276/NSFCU23B2055), Yongjiang Talent Introduction Programme (2022A-238-G), and Fundamental Research Funds for the Central Universities (226-2023-00138). This work was supported by Ant Group.
@@ Appendix B A Decoder Analysis
To explore the effectiveness of our UniHR encoding further, we pair UniHR with different decoders and evaluated them on triple prediction task. In addition to the previously mentioned unified framework UniHR + Transformer, we also experiment on UniHR + ConvE with two scoring strategies. The ConvE (conve) is the decoder customized for triples and its scoring function is $vec\left(\sigma\left(\left[\mathbf{\tilde{h}}_{h};\mathbf{\tilde{e}}_{r}\right]*\psi\right)\right)$, where $\mathbf{\tilde{h}}_{h}$ and $\mathbf{\tilde{e}}_{r}$ represent reshaped 2D embeddings of head entity $h$ and relation $r$, and * is a convolution operator. The $vec\left(\cdot\right)$ and $\psi$ are denoted as the vectorization function and a set of convolution kernels.
@@
Table 6: Base link prediction on FBHE, FBH and DBHE. All baselines’ results are taken from (neste). The best results among all models are written bold, while the second are underlined. The $s_{f}$ and $s_{h}$ denote $\left(f,has\;head\;entity,h\right)\ \left(f,has\;tail\;entity,t\right)$ and $\left(h,r,t\right)$ two types of scoring method respectively.
Due to our special representation, there exists two scoring methods for atomic triples, thus we present the base link prediction results separately for each scoring method. The $s_{f}$ represents scoring triples $\left(f,has\;head\;entity,h\right)\$and $\left(f,has\;tail\;entity,t\right)$, and $s_{t}$ represents scoring $\left(h,r,t\right)$. The performance of base link prediction is shown in Table 6. It can be observed that regardless of the scoring method employed, we both achieve competitive performance, especially with scoring $\left(h,r,t\right)$ on FBH and scoring $\left(f,has\;head\;entity,h\right)\,\left(f,has\;tail\;entity,t\right)$ on DBHE. We attribute the differences in performance under different scoring methods to dataset characteristics. DBHE dataset is relatively smaller, and scoring method $s_{f}$ effectively alleviates overfitting problem. Conversely, for larger datasets FBH, scoring based on $\left(h,r,t\right)$ minimizes information loss.
Table 7 shows the results of triple prediction on three benchmark datasets. Among all baselines, Quate and Bique struggle to model the mapping relationship between atomic facts and nested facts. Furthermore, prior works (bive) do not guarantee that all atomic facts in the nested fact test set are present in the training set as entities, which shifts the problem from a transductive setting to an inductive setting, leading to significant performance gaps between these baselines. On most metrics, our method outperforms BiVE and NestE which are specifically modeled for nested facts. Notably, NestE fully preserves the semantics of atomic facts. However, on the FBHE dataset, UniHR + ConvE achieves an improvement of 0.58 (6.4%) points in MRR and 0.24 (2.4%) points in Hits@10 compared to the state-of-the-art model NestE and the second-best performance after UniHR + Transformer on the FBH and DBHE datasets, demonstrating UniHR’s powerful graph structure encoding capabilities. We also carry out ablation experiments on UniHR + ConvE as shown in Table 7. Performance declines are observed after removing any part of the HiSL module, showing the significance of HiSL for hierarchical encoding.
@@
Table 7: Triple prediction on FBHE, FBH and DBHE.
@@
Appendix C B More Related Works
@@
Most existing techniques in KG representation learning are proposed for triple-based KGs. Among these techniques, knowledge graph embedding (KGE) models (transe; rotate) have received extensive attention due to their effectiveness and simplicity. The idea is to project entities and relations in the KG to low-dimensional vector spaces, utilizing KGE scoring functions to measure the plausibility of triples in the embedding space. Typical methods include TransE (transe), RotatE (rotate), and ConvE (conve).
Depending on the KGE model alone has limitation of capturing complex graph structures, whereas augmenting global structural information with a graph neural network (GNN) (CompGCN; ma-gnn; zhanghe) proves to be an effective approach for enhancement. The paradigm of combining GNN as encoder with KGE scoring function as decoder helps to enhance the performance of KGE scoring function. These GNN methods design elaborate message passing mechanisms to capture the global structural features. Typically, CompGCN (CompGCN) aggregates the joint embedding of entities and relations in the neighborhood via a parameter-efficient way and MA-GNN (ma-gnn) learns global-local structural information based on multi-attention. These methods achieve impressive results on triple-based KGs but are challenging to generalize to beyond-triple KGs. Additionally, these GNNs usually focus more on modeling global information while neglecting local information.
@@
Firstly, real-world data sources naturally contain heterogeneous data forms (ska-bench). For example, Wikidata (wikidata) includes triple-based facts, hyper-relational facts, temporal facts, and even nested facts, financial reports and research papers usually contain various modalities such as tables, images, and text. Unified representation aims to integrate multiple heterogeneous data within a single semantic structure, avoiding the need for cross-type data conversion or independent processing during use. In many real-world scenarios (e.g., question answering, recommendation, or transfer learning), the research community is constantly exploring the benefits of unified representations. For question answering (QA) (rtqa; eog), TrustUQA (trustuqa) demonstrates a unified representation allows the knowledge base to contain a broader range of knowledge types, thereby simplifying the burden of the knowledge retrieval module and enabling the QA system (prgb) to cover a wider variety of questions and answers. For recommendation systems, DURation (unified_rec) with unified representation effectively eliminate information redundancy caused by separate module designs and help capture cross-dimensional relational knowledge, thus improving the comprehensiveness and accuracy of recommendations. Moreover, UniHR not only meets the requirements of real-world KG reasoning tasks (involving multiple types of facts), but we also believe that unified representation itself is beneficial for transfer learning. It not only facilitates transfer learning across different types of KGs, but also enables pre-training across multiple KG types (as demonstrated in Section 5.4). This further promotes more effective pre-training and transfer learning in the field of knowledge graphs.
Appendix D C Detailed Results of Our “Potential”
@@
Appendix E D Dataset Details
@@
Among them, WikiPeople is a dataset derived from Wikidata (wikidata) concerning entities type “human”. WikiPeople filter out the elements which have at least 30 mentions as key-value pairs. WD50K is a high-quality dataset extracting from Wikidata statements and avoiding the potential data leakage which allows triple-based models to memorize main fact in the H-Facts of test set. The “with Q(%)” column in Table 8 denote the number of facts with auxiliary key-value pairs and the “Arity” column denote range of the number of entities in H-Facts.
Nested factual knowledge graph datasets FBH and FBHE (bive) are constructed based on FB15k237 from Freebase (bollacker2008freebase) while DBHE is based on DB15K from DBpedia (DBpedia). FBH contains nested facts that can be only inferred inside the atomic facts, while FBHE and DBHE contain externally-sourced nested relation crawling from Wikipedia articles, e.g., NextAlmaMater and SucceededBy.
Temporal knowledge graph dataset wikidata12K is also a subset of Wikidata (wikidata), which represents the time information $\tau\in\mathcal{T}$ as time intervals.
Appendix F E Hyperparameter Settings